# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ Abstract
General relativity — Tests of general relativity — Gravitational Waves — Ground-based gravitational wave detectors
@@ 1 Introduction
Einstein’s general theory of relativity (GR) stands as the most successful theory of gravity to date. Rigorously tested in weak-field, low-speed, and linear gravity regimes, GR has consistently withstood all scrutiny. Gravitational waves (GWs) are predictions of GR and offer a unique avenue for exploring spacetime dynamics in extreme gravitational conditions. Despite the widespread use of GWs from compact binary coalescences (CBCs) for testing GR, no deviations from the theory have been found so far (Abbott et al., 2016a, 2019a, 2019b, 2021a, 2021b; Yunes et al., 2016; Nair et al., 2019; Silva et al., 2021; Perkins et al., 2021a; Schumacher et al., 2023; Callister et al., 2023; Lagos et al., 2024).
@@
The sensitivity of GW detectors has been continuously improving and LIGO and Virgo detectors are currently witnessing their fourth observing run (O4) with Advanced LIGO and Virgo sensitivity (Aasi et al., 2015) which later will be joined by KAGRA (Akutsu et al., 2021). These detectors will be further upgraded for the fifth observing run (O5) during 2027-2029 (The KAGRA Collaboration, the LIGO Scientific Collaboration, and the Virgo Collaboration, 2022) with A+ sensitivity (Abbott et al., 2018a), and they will eventually be joined by LIGO-India (Iyer et al., 2011; Saleem et al., 2022). Looking further into the future beyond O5, there is a possibility for detectors with A# sensitivity (Fritschel et al., 2022) that are expected to be twice as sensitive as A+. Moreover, there are concrete plans to build next generation (XG) detectors, such as Cosmic Explorer (Evans et al., 2021) and Einstein Telescope (Hild et al., 2011), that are expected to be at least 10 times more sensitive than the current detectors in O4. The first space-borne mission, LISA (Amaro-Seoane et al., 2017), is scheduled to be launched in the mid-2030s, and it might be followed by other missions such as TianQin (Luo et al., 2016; Mei et al., 2021), Taiji (Wu et al., 2021), DECIGO (Kawamura et al., 2006, 2021) and LGWA (Ajith et al., 2024).
@@
With these improvements in sensitivity, thousands of CBCs are expected to be observed with high signal-to-noise ratios (SNRs) (Abbott et al., 2018a). A subset of these mergers will cover extreme regions of the parameter space, including highly spinning and/or strongly precessing binaries, binaries with eccentricity, binaries involving dense matter, etc. Such binaries will have the capability to test GR stringently and constrain beyond-GR effects, if present in the data. For example, higher black hole spins lead to higher curvature outside the horizon (Horowitz et al., 2023), which allows one to place constraints on a variety of higher-derivative or curvature-corrected theories (Jackiw & Pi, 2003; Alexander & Yunes, 2009). More so, the near-horizon region of black holes could potentially access energies as large as the Planck scale that could alter the black hole ringdown spectrum if GR is modified near the event horizon (Barausse et al., 2014; Cardoso et al., 2016a). There is also the possibility that GR may be violated not in the ultraviolet (UV), but rather in the infrared (IR) regime of the theory, aimed at offering an alternative explanation of the dark sector. In this “IR” scenario, extending the reach of GW detectors to lower frequencies may help observe possible deviations from GR in the inspiral phase of CBCs (de Rham & Melville, 2018; ter Haar et al., 2021; Bezares et al., 2022a, 2021).
@@
The majority of tests of GR currently performed rely on waveform models that are compared with the GW data. Often these tests are formulated as null tests where one looks for possible departures from GR by introducing deviation parameters on a given waveform model. No statistically significant deviation from GR has been observed at the level of individual events or for the whole population (Abbott et al., 2021b). However, there were a couple of events in GWTC-3 (Abbott et al., 2023) that suggested GR deviations, though further investigations are needed since these deviations could be due to the use of imperfect waveform models or inadequately understood noise artifacts in the data (Maggio et al., 2023).
@@
Due to the complexity of the physics of compact binary mergers as well as the detector noise modeling, it is extremely important that there is a consensus in the community about the necessary conditions that will warrant a much more comprehensive list of tests to be carried out to vet (or rule out) a potential GR violation claim. There are two aspects to this issue. The first is to identify all possible causes which might lead to a false GR violation. The second is a checklist to be executed upon encountering a strong candidate for GR violation. The objective of this paper is to tackle the first aspect and enumerate an extensive list of scenarios that may appear as violations of GR, when in fact they are not. The second aspect requires us to construct a checklist of items that address other issues such as the statistical significance of the violation, the status of the detector, or if the violation is in contradiction with other experiments or astrophysical observations. A companion paper will address these issues and a possible formulation of a GR violation detection checklist. It is worth noting that a similar effort has been made in Section 7 of Arun et al. (2022), albeit in the context of tests of GR using LISA. Our goal here is to broadly classify different effects that can mimic a GR violation in the context of present- and next-generation ground-based interferometric observational facilities.
@@ 2 Noise Systematics
Current interferometric GW detectors are limited by fundamental noise sources (Aasi et al., 2015) which causes the noise to appear as stationary and Gaussian only over short time scales and ranges of frequency (Abbott et al., 2020a). In reality, however, noise from the detectors is neither Gaussian nor stationary (see, e.g., Abbott et al. (2016b, 2020a); Davis et al. (2021)). It can be relatively easy to spot times of extremely bad data quality in GW data, but the challenge lies with times of subtle data quality issues. The origin of noise sources is notoriously difficult to pinpoint, even for obvious cases of poor data quality. However, it is essential that we understand our noise, remove any bias that noise introduces, and accurately infer the parameters of the observed sources.
@@ 2.1 Non-stationarity
Non-stationarity is a broadband form of noise which causes the statistical properties of the background to change with time. Non-stationarity occurs on the order of tens of seconds in the current LIGO detectors and can be caused by both instrumental and environmental sources (Abbott et al., 2016b; Capote et al., 2024). This form of noise has been shown to affect the estimation of source parameters (Edy et al., 2021; Kumar et al., 2022). Modelled searches typically estimate a detector’s power spectrum over several minutes (Usman et al., 2016; Messick et al., 2017; Venumadhav et al., 2019), which can cause the matched filter to miss the variable nature of the noise, affecting the search sensitivity. One method to account for this is to construct a statistic which tracks the variation of the power spectrum and to normalize the ranking statistic used by the detection pipeline (Venumadhav et al., 2019; Mozzon et al., 2020; Zackay et al., 2021). The method presented in Mozzon et al. (2020) is also used to assess the stationarity of the data around candidate GW events (Davis et al., 2021). This is because non-stationary noise can impact binary neutron star signal parameters (Chatziioannou et al., 2019; Edy et al., 2021) since noise estimates, usually calculated over minutes, fail to capture variations on shorter time scales. As signals from (sufficiently massive) binary black holes are usually shorter than the typical time scale of non-stationary noise, these sources are not thought to be affected.
@@ 2.2 Noise Transients or Glitches
Transient noise artifacts, also known as glitches, are also a common problem in interferometric GW detectors. Glitches can mask or mimic a signal and add to the noise background of transient GW searches (see, e.g., Abbott et al. (2016b, 2018b); Davis et al. (2021)). Glitches occur frequently in all detectors; in the third observing run, the rate of glitches was between 0.29 to 0.32 per minute for LIGO-Hanford, 1.10 to 1.17 per minute for LIGO-Livingston and 0.47 to 1.11 per minute for Virgo (Abbott et al., 2023). The inferred population properties of glitches have been shown to typically exhibit characteristics similar to CBC signals with extreme mass ratios and large spins, compared to the observed astrophysical properties, which tend to have near equal masses and moderate spins (Ashton et al., 2022).
@@
The morphology of glitches, in particular their time duration and the frequency space they affect, can be highly variable between different glitch classes. For example, blip glitches (e.g., Cabero et al. (2019)) are fractions of a second in duration, covering a large bandwidth (e.g., tens to hundreds of Hz) and can mimic a GW signal of high mass compact binaries. We still do not know the origin of these types of glitches as they do not have a known environmental or instrumental coupling, but they appear to have different subcategories that may be caused by different physical mechanisms. In the third observing run, these types of glitches occurred 4 times per hour at LIGO-Livingston and twice per hour at LIGO Hanford (Davis et al., 2021). However, scattering glitches (e.g., Soni et al. (2021)) caused by microseism noise, can be a few seconds long, and present as arches in the time-frequency plane, affecting frequencies below $100$ Hz. These glitches manifest due to a small fraction of laser light scattering off a test mass, hitting a moving surface, and recombining with the main beam. These types of glitches are most prevalent when the ground motion is high. As such they can seriously contaminate hours of data, but not be a concern for weeks at a time.
@@
Tracking the occurrence and emergence of new glitch types can be a challenge. Both LIGO and Virgo take advantage of machine learning frameworks, combined with citizen scientists, to classify glitches based on their morphology in the time-frequency plane. GravitySpy (Zevin et al., 2017) has been in operation since the second observing run, and citizens have helped to classify LIGO glitches into 23 distinct classes (Davis et al., 2021). GWitchHunters (Razzano et al., 2023) helps to classify glitches from the Virgo detector, and has been open to the public since November 2021. Both projects will prove extremely valuable in identifying and understanding glitches in the fourth and future observing runs.
@@
Glitches overlapping or being in the vicinity of a real GW signal can be a huge problem. In fact, in the third observing run 24% of GW events had a glitch within the analysis window for one or more detectors (Abbott et al., 2023). These glitches did not impact the detection of these events, but they had to be mitigated before the source parameters could be accurately estimated. A prime example of this issue first arose in the interpretation of GW170817 where a short glitch occurred 1.1 seconds before the coalescence of the event, lasting only 5 ms (Abbott et al., 2017a). Nonetheless, this noise had to be removed before the parameters of the event could be accurately determined. Macas et al. (2022), for example, shows that certain types of glitches can cause the sky localization to be incorrectly determined for certain types of signals, which can even affect follow-up with large field of view telescopes (i.e., 20 deg2).
@@
There are a number of ways in which noise can be removed or subtracted from the data. Should the noise be broadband in origin then noise subtraction over the course of hours or days is needed. This can be achieved using auxiliary channels which monitor noise sources at different points around an interferometer. A coupling function can then be determined to understand how much a certain type of noise affects the GW channel, and the noise subtracted (Davis et al., 2019; Driggers et al., 2019). This method is optimal when the data are Gaussian and stationary. More recent work has focused on machine learning techniques to cope with data with non-stationary noise couplings (Vajente et al., 2020).
[…]
@@
Glitches will always remain a feature of GW data because as the detector sensitivity improves noise artifacts that were sub-dominant will become more relevant. It is unfeasible to remove them all. New methods are being developed to effectively deduce both source and population parameters by integrating realistic but imperfect data. For example,  Ashton (2023) uses Gaussian processes to replace the traditional GW likelihood. This method, in principle, can model arbitrarily colored noise, non-stationarity, and glitches, to augment the approach to estimate the parameters of sources. In addition, Heinzel et al. (2023) presents a method for inferring the population of GW sources contaminated by blip glitches. They are able to infer the shape parameters of a GW population, whilst simultaneously inferring the population of the glitch background events.
@@
In order to be confident that a signal is indeed a violation of GR, characteristics that may arise due to the noise identified here need to be understood. Work has started in this regard, for example with Kwok et al. (2022). They investigated how an overlapping binary black hole signal with three different glitches can affect tests of GR before and after the glitches were mitigated. Moreover, they only considered a glitch in a single detector out of three and still found a GR deviation when the glitch was not mitigated. The authors also point out that their study is not sufficient to give quantitative statements about the effects of certain glitch classes or mitigation methods on tests of GR. Therefore, their work needs to be extended to assess the amount of GR deviation in different realizations of Gaussian noise, the effect of non-stationarities in the noise background, and the effect of data cleaning methods on mimicking GR deviations.
@@ 2.3 Contamination from Overlapping Signals
As the sensitivity of ground-based GW detectors improves, the chances of observing time-overlapping signals will also increase (Samajdar et al., 2021). This may demand a shift in our detection and parameter estimation strategies since current pipelines, designed for single GW signals, may yield biased results when applied to overlapping signals. However, several studies have shown that the detection (Regimbau et al., 2012; Meacher et al., 2016) and parameter estimation (Samajdar et al., 2021; Pizzati et al., 2022; Relton & Raymond, 2021; Himemoto et al., 2021) of overlapping signals are not a significant concern. Additionally, methods have been proposed to correct biases in cases where overlaps do pose challenges (Antonelli et al., 2021; Janquart et al., 2022; Langendorff et al., 2023).
For example, Regimbau et al. (2012) and Meacher et al. (2016) concluded that the detection of longer signals will not be affected in the presence of multiple signals in data around the same time. More recently, Relton et al. (2022) conducted a more thorough search study with both modeled and unmodelled analyses and put constraints on regimes where the unmodelled searches would perform better when merger times of individual signals are very close to each other. Wu & Nitz (2023) proposed an updated search campaign on overlapping signals where they consider the effects of using the traditional matched filtering and its consequences on estimating the noise properties, as well as the detection rates of overlapping signals. As pointed out in Johnson et al. (2024), the presence of overlapping signals may require us to revisit the definition of the likelihood as well as the assumption that source confusion can be treated as stationary Gaussian noise.
@@
The inference of source parameters is only biased if signals merge very close to each other in the data and differ in SNRs. Possible remedies to this problem have been suggested, either from a Fisher Matrix study (Antonelli et al., 2021) or adapting the signal model accordingly in the Bayesian likelihood (Janquart et al., 2022). Langendorff et al. (2023) used normalizing flows as an avenue to deal with the computational burden coming from multiple-signal analyses in case of overlaps. Moreover, Hu & Veitch (2023) studied the effects of waveform inaccuracy and overlapping signals on tests of GR and concluded that combining signals can lead to false GR deviations in case of multiple signal overlaps. More recently, Dang et al. (2023) extended this study to higher post-Newtonian (PN) deformation parameters. They concluded that although a non-negligible number of overlapping signals can lead to false GR violations at the individual event level, when the results are combined, the biases tend to smoothen out, leading to a preference for GR at the population level inference (We discuss the effects of population-level analyses on tests of GR in more detail in Section 4.4.)
@@
All these studies focussed on overlaps arising in the data of XG detectors, since the probability of observing overlapping signals in the era of A+ sensitivity (Abbott et al., 2018a) or Voyager (Adhikari et al., 2020) is very small (Samajdar et al., 2021). However, it is likely that a quiet GW signal below the detection threshold is present along with the dominant GW signal in the data (Abbott et al., 2024). This will not pose a problem for estimating individual source parameters, but issues may arise when combining multiple signals, where sub-threshold events collectively act as background or confusion noise (Reali et al., 2022, 2023). Although Reali et al. (2022, 2023) considered signals in the XG era only, we might need to consider the effect of a confusion-noise-like background in O5 or A# era in the context of testing GR. Moreover, quieter signals may result in imperfect subtraction of the GW model from data when following the definition of likelihood to infer source properties under the assumption of stationary, Gaussian noise. Consequently, combining results across multiple signals to infer population properties could gradually accumulate biases from each single-signal analysis, potentially mimicking noise properties (Johnson et al., 2024) and introducing deviations from GR.

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Phys. Comm. Rep.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
Einstein’s general theory of relativity (GR) stands as the most successful theory of gravity to date. Rigorously tested in weak-field, low-speed, and linear gravity regimes, GR has consistently withstood all scrutiny. Gravitational waves (GWs) are predictions of GR and offer a unique avenue for exploring spacetime dynamics in extreme gravitational conditions. Despite the widespread use of GWs from compact binary coalescences (CBCs) for testing GR, no deviations from the theory have been found so far (e.g., [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]).
@@
The sensitivity of GW detectors has been continuously improving and LIGO and Virgo detectors are currently witnessing their fourth observing run (O4) with Advanced LIGO and Virgo sensitivity [13] which later will be joined by KAGRA [14]. These detectors will be further upgraded for the fifth observing run (O5) during 2027-2029 [15] with A+ sensitivity [16], and they will eventually be joined by LIGO-India [17, 18]. Looking further into the future beyond O5, there is a possibility for detectors with A# sensitivity [19] that are expected to be twice as sensitive as A+. Moreover, there are concrete plans to build next generation (XG) detectors, such as Cosmic Explorer [20] and Einstein Telescope [21], that are expected to be at least 10 times more sensitive than the current detectors in O4. The first space-borne mission, LISA [22], is scheduled to be launched in the mid-2030s, and it might be followed by other missions such as TianQin [23, 24], Taiji [25], DECIGO [26, 27] and LGWA [28].
@@
With these improvements in sensitivity, thousands of CBCs are expected to be observed with high signal-to-noise ratios (SNRs) [16]. A subset of these mergers will cover extreme regions of the parameter space, including highly spinning and/or strongly precessing binaries, binaries with eccentricity, binaries involving dense matter, etc. Such binaries will have the capability to test GR stringently and constrain beyond-GR effects, if present in the data. For example, higher black hole spins lead to higher curvature outside the horizon [29], which allows one to place constraints on a variety of higher-derivative or curvature-corrected theories [30, 31]. More so, the near-horizon region of black holes could potentially access energies as large as the Planck scale that could alter the black hole ringdown spectrum if GR is modified near the event horizon [32, 33]. There is also the possibility that GR may be violated not in the ultraviolet (UV), but rather in the infrared (IR) regime of the theory, aimed at offering an alternative explanation of the dark sector. In this “IR” scenario, extending the reach of GW detectors to lower frequencies may help observe possible deviations from GR in the inspiral phase of CBCs [34, 35, 36, 37].
@@
The majority of tests of GR currently performed rely on waveform models that are compared with the GW data. Often these tests are formulated as null tests where one looks for possible departures from GR by introducing deviation parameters on a given waveform model. No statistically significant deviation from GR has been observed at the level of individual events or for the whole population [5]. However, there were a couple of events in GWTC-3 [38] that suggested GR deviations, though further investigations are needed since these deviations could be due to the use of imperfect waveform models or inadequately understood noise artifacts in the data [39].
@@
Due to the complexity of the physics of compact binary mergers as well as the detector noise modeling, it is extremely important that there is a consensus in the community about the necessary conditions that will warrant a much more comprehensive list of tests to be carried out to vet (or rule out) a potential GR violation claim. There are two aspects to this issue. The first is to identify all possible causes which might lead to a false GR violation. The second is a checklist to be executed upon encountering a strong candidate for GR violation. The objective of this paper is to tackle the first aspect and enumerate an extensive list of scenarios that may appear as violations of GR, when in fact they are not. The second aspect requires us to construct a checklist of items that address other issues such as the statistical significance of the violation, the status of the detector, or if the violation is in contradiction with other experiments or astrophysical observations. A companion paper will address these issues and a possible formulation of a GR violation detection checklist. It is worth noting that a similar effort has been made in Section 7 of [40], albeit in the context of tests of GR using LISA. Our goal here is to broadly classify different effects that can mimic a GR violation in the context of present- and next-generation ground-based interferometric observational facilities.
@@ 2 Noise Systematics
Current interferometric GW detectors are limited by fundamental noise sources [13] which causes the noise to appear as stationary and Gaussian only over short time scales and ranges of frequency [41]. In reality, however, noise from the detectors is neither Gaussian nor stationary (see, e.g., [42, 41, 43]). It can be relatively easy to spot times of extremely bad data quality in GW data, but the challenge lies with times of subtle data quality issues. The origin of noise sources is notoriously difficult to pinpoint, even for obvious cases of poor data quality. However, it is essential that we understand our noise, remove any bias that noise introduces, and accurately infer the parameters of the observed sources.
@@ 2.1 Non-stationarity
Non-stationarity is a broadband form of noise which causes the statistical properties of the background to change with time. Non-stationarity occurs on the order of tens of seconds in the current LIGO detectors and can be caused by both instrumental and environmental sources [42, 44], such as detector’s lockloss [45], variable seismic motion, thunderstorms [43], magnetic effects from lightning strokes [46], and radio frequency interference [42, 43]. This form of noise has been shown to affect the estimation of source parameters [47, 48]. Modelled searches typically estimate a detector’s power spectrum over several minutes [49, 50, 51], which can cause the matched filter to miss the variable nature of the noise, affecting the search sensitivity. One method to account for this is to construct a statistic which tracks the variation of the power spectrum and to normalize the ranking statistic used by the detection pipeline [51, 52, 53]. The method presented in [52] is also used to assess the stationarity of the data around candidate GW events [43]. This is because non-stationary noise can impact binary neutron star signal parameters [54, 55] since noise estimates, usually calculated over minutes, fail to capture variations on shorter time scales. As signals from sufficiently massive binary black holes are usually shorter than the typical time scale of non-stationary noise, these sources are not thought to be affected. However, due to their long duration in the sensitivity band, sub-solar mass binary black hole searches will be affected, especially in the XG era where the signal duration could be up to several days.
@@ 2.2 Noise Transients or Glitches
Transient noise artifacts, also known as glitches, are also a common problem in interferometric GW detectors. Glitches can mask or mimic a signal and add to the noise background of transient GW searches (see, e.g., [42, 56, 43]). Glitches occur frequently in all detectors; in the third observing run, the rate of glitches was between 0.29 to 0.32 per minute for LIGO-Hanford, 1.10 to 1.17 per minute for LIGO-Livingston and 0.47 to 1.11 per minute for Virgo [38]. The inferred population properties of glitches have been shown to typically exhibit characteristics similar to CBC signals with large mass ratios and large spins, in contrast to the observed astrophysical properties, which tend to have near equal masses and moderate spins [57]. This is because CBC signals with large mass ratios and large spins can have more ‘irregular’ waveform morphologies compared to equal-mass, non-spinning CBCs from the twisting-up effects due to precession. Therefore, this class of signals has a better chance of fitting well with the terrestrial disturbance produced by glitches that lack a CBC signal’s typical chirping-up characteristics.
@@
The morphology of glitches, in particular their time duration and the frequency space they affect, can be highly variable between different glitch classes. For example, blip glitches (e.g., [58]) are fractions of a second in duration, covering a large bandwidth (e.g., tens to hundreds of Hz) and can mimic a GW signal of high mass compact binaries. We still do not know the origin of these types of glitches as they do not have a known environmental or instrumental coupling, but they appear to have different subcategories that may be caused by different physical mechanisms. In the third observing run, these types of glitches occurred 4 times per hour at LIGO-Livingston and twice per hour at LIGO Hanford [43]. However, scattering glitches (e.g., [59]) caused by microseism noise, can be a few seconds long, and present as arches in the time-frequency plane, affecting frequencies below $100$ Hz. These glitches manifest due to a small fraction of laser light scattering off a test mass, hitting a moving surface, and recombining with the main beam. These types of glitches are most prevalent when the ground motion is high. As such they can seriously contaminate hours of data, but not be a concern for weeks at a time.
@@
Tracking the occurrence and emergence of new glitch types can be a challenge. Both LIGO and Virgo take advantage of machine learning frameworks, combined with citizen scientists, to classify glitches based on their morphology in the time-frequency plane. GravitySpy [60] has been in operation since the second observing run, and citizens have helped to classify LIGO glitches into 23 distinct classes [43]. GWitchHunters [61] helps to classify glitches from the Virgo detector, and has been open to the public since November 2021. Both projects will prove extremely valuable in identifying and understanding glitches in the fourth and future observing runs.
@@
Glitches overlapping or being in the vicinity of a real GW signal can be a huge problem. In fact, in the third observing run 24% of GW events had a glitch within the analysis window for one or more detectors [38]. These glitches did not impact the detection of these events, but they had to be mitigated before the source parameters could be accurately estimated. A prime example of this issue first arose in the interpretation of GW170817 where a short glitch occurred 1.1 seconds before the coalescence of the event, lasting only 5 ms [62]. Nonetheless, this noise had to be removed before the parameters of the event could be accurately determined. Macas et al. [63], for example, shows that certain types of glitches can cause the sky localization to be incorrectly determined for certain types of signals, which can even affect follow-up with large field of view telescopes (i.e., 20 deg2).
@@
There are a number of ways in which noise can be removed or subtracted from the data. Should the noise be broadband in origin then noise subtraction over the course of hours or days is needed. This can be achieved using auxiliary channels which monitor noise sources at different points around an interferometer. A coupling function can then be determined to understand how much a certain type of noise affects the GW channel, and the noise subtracted [64, 65]. This method is optimal when the data are Gaussian and stationary. More recent work has focused on machine learning techniques to cope with data with non-stationary noise couplings [66].
For short instances of transient noise that may be in the vicinity of an event, there are a few methods which are currently used. A window function can be applied to zero out the glitch; this method is known as gating [49, 67]. Gating has the benefit of being quick, however uncontaminated data will also be removed using this method, as the window function needs to be smoothly applied to avoid adding filtering artifacts to the data. Hence, this method is not appropriate if the glitch is not well localized in time and is close to an event’s coalescence time. A more robust method is to model a glitch with a time-frequency wavelet reconstruction and use this to subtract it from the data; this method is applied using the BayesWave algorithm [68]. This method has been used to great effect in the third observing run [38]. Another method, called gwsubtract, uses data from an auxiliary witness to the noise to subtract the noise from the GW channel [64, 69]. This was done for the first time around the event GW200129 [38], which seems to exhibit characteristics consistent with spin induced orbital precession [70]. However, Payne et al. [71] find that residual data quality issues leftover from this cleaning process may be the origin of the precession observed in GW200129. Moreover, in a ringdown analysis of GW200129 [39] found a deviation from GR in the peak of the GW amplitude while employing a nonprecessing SEOBNRv4HM_PA model [72, 73, 74] but they ascribe it to waveform systematics (modeling of spin precession) or data-quality issues (glitch mitigation procedures). Regardless, this example of GW200129 highlights the complexities and care that need to be taken when removing glitches from GW data and interpreting results from inference analyses.
@@
Glitches will always remain a feature of GW data because as the detector sensitivity improves noise artifacts that were sub-dominant will become more relevant. It is unfeasible to remove them all. New methods are being developed to effectively deduce both source and population parameters by integrating realistic but imperfect data. For example, Ashton et al. [75] uses Gaussian processes to replace the traditional GW likelihood. This method, in principle, can model arbitrarily colored noise, non-stationarity, and glitches, to augment the approach to estimate the parameters of sources. In addition, Heinzel et al.[76] presents a method for inferring the population of GW sources contaminated by blip glitches. They are able to infer the shape parameters of a GW population, whilst simultaneously inferring the population of the glitch background events.
@@
In order to be confident that a signal is indeed a violation of GR, characteristics that may arise due to the noise identified here need to be understood. Work has started in this regard, for example with [77]. They investigated how an overlapping binary black hole signal with three different glitches can affect tests of GR before and after the glitches were mitigated. Moreover, they only considered a glitch in a single detector out of three and still found a GR deviation when the glitch was not mitigated. The authors also point out that their study is not sufficient to give quantitative statements about the effects of certain glitch classes or mitigation methods on tests of GR. Therefore, their work needs to be extended to assess the amount of GR deviation in different realizations of Gaussian noise, the effect of non-stationarities in the noise background, and the effect of data cleaning methods on mimicking GR deviations.
@@ 2.3 Contamination from Overlapping Signals
As the sensitivity of ground-based GW detectors improves, the chances of observing time-overlapping signals will also increase [78]. This may demand a shift in our detection and parameter estimation strategies since current pipelines, designed for single GW signals, may yield biased results when applied to overlapping signals. However, several studies have shown that the detection [79, 80] and parameter estimation [78, 81, 82, 83] of overlapping signals are not a significant concern. For example, [79] and [80] showed that it is possible to detect and discern overlapping signals from binary neutron stars using a matched filtering algorithm. Relton et al. [84] conducted a more thorough study with both modeled and unmodelled search analyses and found that both analyses can detect overlapping signals from binary black holes when merging $>1$ s apart. However, unmodelled analysis can identify overlapping signals merging within $<1$ s while modeled analysis can only identify only one of the two overlapping signals. Himemoto et al. [83] thoroughly explored the parameter space and concluded that overlapping signals do not lead to large biases in parameter estimation provided the coalescence times and redshifted chirp masses of the two overlapping signals differ by at least $10^{-2}$ s and $10^{-4}M_{\odot}$ for binary neutron star mergers and $10^{-1}$ s and $10^{-1}M_{\odot}$ for binary black hole mergers, respectively.
Nonetheless, overlapping signals do pose biases in both detection and parameter estimation of sources and methods have been proposed to correct those biases [85, 86, 87, 88]. For example, Wu & Nitz [85] pointed out that overlapping signals reduce the search sensitivity by changing the noise’s amplitude spectral density and proposed an updated search campaign on overlapping signals using the single-detector signal subtraction method. Johnson et al. [89] pointed out that the presence of overlapping signals may require us to revisit the definition of the likelihood as well as the assumption that source confusion can be treated as stationary Gaussian noise. Possible remedies to the bias in source parameter inference have been suggested, either from a Fisher Matrix study [86] or adapting the signal model accordingly in the Bayesian likelihood [87]. Langendorff et al. [88] used normalizing flows as an avenue to deal with the computational burden coming from multiple-signal analyses in case of overlaps.
@@
Moreover, Hu & Veitch [90] studied the effects of waveform inaccuracy and overlapping signals on tests of GR and demonstrated that when combining results from multiple signals with overlaps, the deviation from GR decreases when waveform models are perfect (no waveform inaccuracy), but inaccurate waveform modeling can lead to a false deviation of GR for overlapping signals. Dang et al. [91] extended this study to higher post-Newtonian (PN) deformation parameters. They concluded that although a non-negligible number of overlapping signals can lead to false GR violations at the individual event level, when the results are combined, the biases tend to smoothen out, leading to a preference for GR at the population level inference. (We discuss the effects of population-level analyses on tests of GR in more detail in Section 4.4.)
@@
All these studies focussed on overlaps arising in the data of XG detectors, since the probability of observing overlapping signals in the era of A+ sensitivity [16] or Voyager [92] is very small [78]. However, it is likely that a quiet GW signal below the detection threshold is present along with the dominant GW signal in the data [93]. This will not pose a problem for estimating individual source parameters, but issues may arise when combining multiple signals, where sub-threshold events collectively act as background or confusion noise [94, 95]. Although [94, 95] considered signals in the XG era only, we might need to consider the effect of a confusion-noise-like background in O5 or A# era in the context of testing GR. Moreover, quieter signals may result in imperfect subtraction of the GW model from data when following the definition of likelihood to infer source properties under the assumption of stationary, Gaussian noise. Consequently, combining results across multiple signals to infer population properties could gradually accumulate biases from each single-signal analysis, potentially mimicking noise properties [89] and introducing deviations from GR.